# One Tool Is Enough: Reinforcement Learning for Repository-Level LLM Agents

### v4

@@ Abstract
Locating the files and functions requiring modification in large open-source software (OSS) repositories is challenging due to their scale and structural complexity. Existing large language model (LLM)-based methods typically treat this as a repository-level retrieval task and rely on multiple auxiliary tools, which overlook code execution logic and complicate model control. We propose RepoNavigator, an LLM agent equipped with a single execution-aware tool—jumping to the definition of a invoked symbol. This unified design reflects the actual flow of code execution while simplifying tool manipulation. RepoNavigator is trained end-to-end via Reinforcement Learning (RL) directly from a pretrained model, without any closed-source distillation. Experiments demonstrate that RL-trained RepoNavigator achieves state-of-the-art performance, with the 7B model outperforming 14B baselines, the 14B model surpassing 32B competitors, and even the 32B model exceeding closed-source models such as Claude-3.7. These results confirm that integrating a single, structurally grounded tool with RL training provides an efficient and scalable solution for repository-level issue localization.
@@ 1 Introduction
With the rapid advancement of Large Language Models (LLMs) (Liu et al., 2024; Team, 2024; Yang et al., 2025a), equipping LLMs with pre-built tools to form LLM agents has become a common paradigm for expanding their capabilities (Shen, 2024; Yuan et al., 2024; Lu et al., 2024). In the domain of software engineering (SWE), although LLM agents can effectively handle simple programming tasks (Hui et al., 2024; Guo et al., 2024a), their ability to operate on large-scale open-source software (OSS) repositories remains limited. SWE-bench (Jimenez et al., 2023) currently serves as the most comprehensive benchmark for evaluating whether LLMs can resolve real-world GitHub issues. All pretrained LLMs can not process the whole repository directly due to context limits. While SWE-Agent (Jimenez et al., 2023) provides moderate gains, it remains far from enabling robust repository-level reasoning.
@@
Most existing agents rely on test-time scaling applied directly to pretrained LLMs (Liu et al., 2023; Chen et al., 2025; Schmidgall et al., 2025). In software engineering (SWE) tasks, tools are essential rather than optional: real-world repositories are far larger than the context window of current LLMs, making it impossible to process an entire codebase in a single forward pass. Agents must therefore iteratively invoke tools to retrieve partial information from the repository and interleave natural-language reasoning with tool calls.
@@
However, mainstream LLM agents (Chen et al., 2025; Liu et al., 2025; Xiang et al., 2025; Wang et al., 2023; Chen et al., 2024) are rarely exposed to such agentic interaction patterns during pretraining and typically acquire tool usage only through few-shot prompting. Such in-context demonstrations are insufficient for learning complex multi-step tool-chaining behaviors, especially under limited context windows. Moreover, because tool definition spaces are effectively unbounded, pretrained models cannot fully internalize their semantics without post-training. To mitigate these issues, post-training paradigms such as Supervised Finetuning (SFT) (Ma et al., 2025) and Reinforcement Learning with Verifiable Rewards (RLVR) (Yu et al., 2025a; Yue et al., 2025) have been applied, with promising results in domains including retrieval agents (Jin et al., 2025), GUI agents (Hong et al., 2024), and math agents (Yan et al., 2025).
@@
Directly training an agent to fix software issues, however, remains difficult. A single bug often admits multiple valid patches, making string-level evaluation unreliable. The only precise evaluation method requires executing candidate patches inside a dedicated Docker environment for each repository (Luo et al., 2025), which is prohibitively expensive. To make training more tractable, we adopt a simplified yet widely generalizable assignment: issue localization. Prior work shows that a software issue becomes substantially easier to resolve once the relevant functions and files are correctly identified (Chen et al., 2025; Ma et al., 2025; Xia et al., 2024; Jiang et al., 2025). Since modern OSS repositories contain a significant amount of code—far beyond any LLM’s context window—localization drastically reduces the search space and improves downstream solvability. Crucially, localization outputs a discrete set of paths, enabling verifiable, string-level evaluation that is compatible with scalable training frameworks such as SFT and RLVR.
@@
Existing localization agents (Ma et al., 2025; Chen et al., 2025; He et al., 2025) typically rely on multiple tools, including SearchClass, SearchMethods, and GetImports. Although effective to some extent, these tools considers high-level abstractions (classes, function, etc) of programing languages, which do not reflect how code actually executes. High-level abstractions, such as classes or inheritance, disappear after compilation, leaving only sequential execution and jump operations. Since modern LLMs already excel at modeling sequential dependencies, we focus on enhancing their ability to jump across the repository—that is, to follow and inspect the source definition of symbols as they appear in execution. To this end, we introduce a single, structurally grounded tool: jump, which retrieves the precise definition of a given symbol. Details of this tool are provided in Sec. 3.3.
@@
Our main contributions are threefold: (1) We propose the first repo-level localization agent trained on reinforcement learning directly from the pretrained model, regardless of distillation from a close-source model. (2) We design a repository-navigation agent that operates by performing realistic jump operations aligned with actual execution semantics. (3) We demonstrate that one unified tool significantly improves efficiency and controllability compared to multi-tool pipelines.
@@ 2.1 Agentic Training
LLM agents are promising methods to equip models with complex tools while reasoning (Li et al., 2024; Huang et al., 2024; Guo et al., 2024b). However, because most pretrained LLMs are trained on texts only and developers can define any tools, most tools are out-of-domain (OOD) for LLMs. Even for the most powerful models, failures often happen when calling the new-defined tools due to wrong calling format or failed parameter parsing. Thus, training a LLM to master new-defined tool is critical for LLM agents. Intuitively, the tool-calling trajectories can be generated by a more powerful LLM, and such trajectories can be used to train a student model via supervised finetuning (SFT) (Chen et al., 2025). However, this pipeline requires a stronger teacher model which has capability to master the tool. Recently, more methods have emerged with no teacher-model required. Rejected-sampled finetuning (RFT) (Ahn et al., 2024) utilizes generated trajectories of the agent itself via multiple rollouts. Agentic RL (Jin et al., 2025) is an on-policy RLVR methods requiring only the result for verifiying trajectories. Such training methods yield remarkable results when the tools are search engines (Jin et al., 2025), python executer (Jimenez et al., 2023), calculator (Yan et al., 2025), and visual models (Gupta and Kembhavi, 2023).
Figure 2: Overview of our RepoNavigator. During the rollout phrase, the agent can invoke the jump tool, and the language server will return the definition code of the symbol. This process is trained by reinforcement learning.
@@ 2.2 Software Engineering Agents
The introduction of SWE-bench (Jimenez et al., 2023; Yang et al., 2024b) has motivated a range of agentic pipelines for software engineering (SWE) tasks. Among them, SWE-agent (Yang et al., 2024a) and OpenHands (Wang et al., 2025a) are widely adopted frameworks that equip agents with tools for interacting with computing environments (terminal in Linux). Workflow-based methods such as Agentless (Xia et al., 2024) decomposes issue resolution into localization, repair, and validation subproblems. PatchPilot (Li et al., 2025) adds reproduction and refinement process in the workflow. CORTEXA (Sohrabizadeh et al., ) reflects the repository into an embedding space to enable efficient retrieval. Chen et al. (2025) builds the respository as a graph and applied graph-level searching tools for localization, and Wang et al. (2025a) furthermore integrated commit history as agent memory. RepoLens (Wang et al., 2025b) equip conceptual information of the respository to enable repo-level understanding. These pipelines are training-free, and yield competitive results.
To enable task-specific training, DeepSWE (Luo et al., 2025) and SWE-Swiss (He et al., 2025) employ reinforcement learning and achieve strong performance. SWE-Gym (Pan et al., 2024) builds a virtual environment to train an agent for this task. However, end-to-end training remains costly because patch evaluation requires executing Docker containers across numerous repositories, and the Docker container hinders guided search methods due to its non-serializablity (Zainullina et al., 2025). Consequently, issue localization has emerged as a computationally efficient alternative, aiming to identify faulty components—at file or function level—rather than generating full patches.
@@ 3.3 Jump: Symbol Resolution
For a name symbol xx, candidate definitions are searched along a scope chain
@@
. For attribute symbols, it computes a (possibly union-valued) type tt for the receiver expression aa using type annotations, assignment flow analysis, function return types, and stub files (.pyi). Member resolution is then defined as
@@
3.4 Reasoning–Action Loop
@@ 3.5 Reinforcement Learning
Table 1: Comparison of different agent pipelines on function-level and file-level metrics. We use Qwen2.5-Instruct series as our base model. Bold numbers denote the best performance among same-size models; underline numbers denote the best training-free performance among same-size models; yellow background illustrates training-free RepoNavigator; blue background illustrates RepoNavigator trained with GRPO.
@@
Close-source Models
@@ 4.1 Experimnent Setup
We extract valid samples from SWE-smith (Yang et al., 2025b) to form the training set. We apply Qwen2.5-7B-Instruct with RepoNavigator to sample each data for 16 times. A sample is abandoned if all 16 scores are zero. For validation, we test our method on SWE-bench-verified (Jimenez et al., 2023), which is a human-verified subset of SWE-bench. We additionally test our method on a subset of SWE-bench-pro (Yang et al., 2025b) (which is a new and more difficult benchmark) for generalization. For ground-truth locations, we directly use the locations in golden patches. All datasets are open-source and are built on real-world github issues.
@@
CoSIL Training Free 8.64 3.33 4.58 2.87 26.64 8.47 12.11 7.70
Agentless Training Free 12.82 6.94 8.05 5.73 39.41 13.15 18.89 12.35
@@
CoSIL Training Free 10.73 4.67 5.96 3.94 34.31 9.97 14.81 9.30
Agentless Training Free 10.49 6.75 7.41 5.28 41.42 13.42 19.02 12.37
@@
CoSIL Training Free 15.00 6.35 8.14 5.21 45.37 13.04 19.42 12.36
Agentless Training Free 11.08 7.31 7.98 5.80 43.07 13.89 20.07 13.11
RepoSearcher Training Free 2.00 1.29 1.45 1.00 13.51 3.43 5.31 3.24
@@
For the 7B model, we conduct GRPO with 8 Tesla-A100-80G GPUs. For the 14B and 32B model, we train it with 16 Tesla-A100-80G GPUs. We apply verl (Shen, 2024) as the training framework, and we apply vLLM (Kwon et al., 2023) as the inference engine. We train the model for 1 epoch, while the training batch size is fixed to 128 on 4k training samples filtered from SWE-smith, with maximum prompt length and max response length both set to 10240. Additionally, we rollout 8 times for each sample, and the temperature is set to 1.0 to encourage exploration. We use greedy decoding in the inference stage to ensure stable performance. More implementation details are provided in Appendix. B.
@@ 4.2 Effectiveness
We compare our method against Locagent (Chen et al., 2025), CoSIL (Jiang et al., 2025), Agentless (Xia et al., 2024), Orcaloca (Yu et al., 2025b), and RepoSearcher (Ma et al., 2025). Detailed explaination of baseline methods are presented in Appendix. A.
@@ Results
As illustrated in Table. 1, on balanced metrics (Sample-F1 and IoU) for both function-level and file-level localization, our method surpasses all baseline methods with the same model size. Moreover, if we train RepoNavigator with GRPO, the 7B model surpasses 14B baselines, and our 14B model surpasses 32B baselines on Sample-F1 and IoU. This contributes to the validness of RepoNavigator furthermore. Although some baselines have a higher recall score, they are prompted to generate a fixed number of results, resulting in low precision, sample-F1 and IoU. This indicates that RepoNavigator behaves more conservatively and generates less wrong locations. For 14B and 32B models, RepoNavigator achieves SOTA among all training-free methods. This implies that the tool we implement is effective and promising, and our single tool pipeline is better than previous multiple tools pipelines.
@@
To assess the generalizability of RepoNavigator, we present its performance on Python samples from the SWE-bench-Pro dataset (Yang et al., 2025b) in Table 2. The results on this dataset are consistent with those observed on SWE-bench_Verified. While we cannot fully exclude the potential influence of data leakage in SWE-bench_Verified, we can make a stronger claim regarding SWE-bench_Pro, as it was released after the publication of the Qwen2.5 series.
Figure 3: Ablation study: comparison between RepoNavigator with training free, RFT, GRPO with pure outcome and hybrid reward on Qwen2.5-7B-Instruct.
@@ 4.3 Training Strategy Comparison
To explore the capability of GRPO on agentic training, we compare GRPO against RFT-only and RFT+GRPO. As presented in Fig. 3, directly training with GRPO outperformes RFT-only and RFT+GRPO. Moreover, although RFT has accetable performance, the more steps RFT proceeds, the less improvement GRPO makes after the cold start. This conclusion contradicts with previous SWE agents trained with RL (Ma et al., 2025), however, it aligns with the broader field of reinforcement learning, where RFT and SFT (as a cold start) is effective only when the pretrained model is not strong enough (Guo et al., 2024a). When the pretrained model is strong enough and data is high-quality, directly training a model with RL is better than training after SFT (RFT) as its cold start.

### v5

# One Tool Is Enough: Reinforcement Learning of LLM Agents for Repository-Level Code Navigation

@@ Abstract
Locating files and functions requiring modification in large software repositories is challenging due to their scale and structural complexity. Existing LLM-based methods typically treat this as a repository-level retrieval task and rely on multiple auxiliary tools, which often overlook code execution logic and complicate model control. We propose RepoNavigator, an LLM agent equipped with a single execution-aware tool: jumping to the definition of an invoked symbol. This unified design reflects the actual flow of code execution while simplifying tool manipulation. RepoNavigator is trained end-to-end via Reinforcement Learning (RL) directly from a base pretrained model, without relying on closed-source distillation. Experiments demonstrate that RL-trained RepoNavigator achieves state-of-the-art performance, with the 7B model outperforming 14B baselines, the 14B model surpassing 32B competitors, and the 32B model exceeding closed-source models such as GPT-5 on most metrics. These results confirm that integrating a single, structurally grounded tool with RL training provides an efficient and scalable solution for repository-level issue localization.
@@ 1 Introduction
With the rapid advancement of Large Language Models (LLMs) (Liu et al., 2024a; Team, 2024; Yang et al., 2025a), equipping LLMs with pre-built tools to form LLM agents has become a common paradigm for expanding their capabilities (Shen, 2024; Yuan et al., 2024; Lu et al., 2024). In the domain of software engineering (SWE), although LLM agents can effectively handle simple programming tasks (Hui et al., 2024; Guo et al., 2024a), their ability to operate on large-scale software repositories remains limited. SWE-bench (Jimenez et al., 2023) currently serves as the most comprehensive benchmark for evaluating whether LLMs can resolve real-world GitHub issues. All pretrained LLMs cannot process the whole repository directly due to context limits (Yang et al., 2024b). While SWE-Agent (Jimenez et al., 2023) provides moderate gains, it remains far from enabling robust repository-level reasoning.
@@
Most existing agents rely on test-time scaling applied directly to pretrained LLMs (Liu et al., 2023; Chen et al., 2025; Schmidgall et al., 2025). In SWE tasks, tools are essential rather than optional: real-world repositories are far larger than the context window of current LLMs, making it impossible to process an entire codebase in a single forward pass. Agents must therefore iteratively invoke tools to retrieve partial information from the repository and interleave natural-language reasoning with tool calls.
@@
Mainstream LLM agents (Chen et al., 2025; Liu et al., 2025; Xiang et al., 2025; Wang et al., 2023; Chen et al., 2024) are rarely exposed to such agentic interaction patterns during pretraining and typically acquire tool usage only through few-shot prompting which is insufficient for learning complex multi-step tool-chaining behaviors. Moreover, because tool definition spaces are effectively unbounded, pretrained models cannot fully internalize their semantics without post-training. To mitigate these issues, post-training paradigms such as Supervised Finetuning (SFT) (Ma et al., 2025) and Reinforcement Learning with Verifiable Rewards (RLVR) (Yu et al., 2025a; Yue et al., 2025) have been applied, with promising results in domains including retrieval agents (Jin et al., 2025a), GUI agents (Hong et al., 2024), and math agents (Yan et al., 2025).
@@
Directly training an agent to fix software issues, however, remains difficult. A single bug often admits multiple valid patches, making string-level evaluation unreliable. The only precise evaluation method requires executing candidate patches inside a dedicated Docker environment for each repository (Luo et al., 2025), which is prohibitively expensive CPU resource for supervised training. To make training more tractable, we adopt a simplified yet widely generalizable assignment: issue localization. Prior work shows that a software issue becomes substantially easier to resolve once the relevant functions and files are correctly identified (Chen et al., 2025; Ma et al., 2025; Xia et al., 2024; Jiang et al., 2025). Since modern software repositories contain a significant amount of code—far beyond any LLM’s context window—localization drastically reduces the search space and improves downstream solvability. Crucially, localization outputs a discrete set of paths, enabling verifiable, string-level evaluation that is compatible with scalable training frameworks such as SFT and RLVR (string-level evaluation requires little resource).
@@
Existing localization agents (Ma et al., 2025; Chen et al., 2025; He et al., 2025) typically rely on multiple tools, including SearchClass, SearchMethods, and GetImports. Although effective to some extent, these tools considers high-level abstractions (classes, function, etc) of programming languages, which do not reflect how code actually executes. High-level abstractions, such as classes or inheritance, disappear after compilation, leaving only sequential execution and jump operations. Since modern LLMs already excel at modeling sequential dependencies, we focus on enhancing their ability to jump across the repository—that is, to follow and inspect the source definition of symbols as they appear in execution. To this end, we introduce a single, structurally grounded tool: jump, which retrieves the precise definition of a given symbol. Our agent invokes this single tool to navigate in the repository, as depicted in Fig. 1. Details of this tool are provided in Sec. 3.3.
@@
Our main contributions are threefold: (1) We propose the first repo-level localization agent trained on reinforcement learning directly from the pretrained model, regardless of distillation from a closed-source model. (2) We design a repository-navigation agent that operates by performing realistic jump operations aligned with actual execution semantics. (3) We demonstrate that one unified tool significantly improves efficiency and controllability compared to multi-tool pipelines.
@@ 2.1 Agentic Training
LLM agents are promising methods to equip models with complex tools while reasoning (Li et al., 2024; Huang et al., 2024; Guo et al., 2024b). However, because most pretrained LLMs are trained on texts only and developers can define any tools, most tools are out-of-domain (OOD) for LLMs. Even for the most powerful models, failures often happen when calling the new-defined tools due to wrong calling format or failed parameter parsing. Thus, training a LLM to master new-defined tool is critical for LLM agents. Intuitively, the tool-calling trajectories can be generated by a more powerful LLM, and such trajectories can be used to train a student model via supervised finetuning (SFT) (Chen et al., 2025). However, this pipeline requires a stronger teacher model which has capability to master the tool. Recently, more methods have emerged with no teacher-model required. Rejected-sampled finetuning (RFT) (Ahn et al., 2024) utilizes generated trajectories of the agent itself via multiple rollouts. Agentic RL (Jin et al., 2025a) is an on-policy RLVR methods requiring only the result for verifying trajectories. Such training methods yield remarkable results when the tools are search engines (Jin et al., 2025a), python executer (Jimenez et al., 2023), calculator (Yan et al., 2025), and visual models (Gupta and Kembhavi, 2023).
Figure 2: Overview of our RepoNavigator. During the rollout phase, the agent can invoke the jump tool, and the language server will return the definition code of the symbol. This process is trained by reinforcement learning.
@@ 2.2 Software Engineering Agents
The introduction of SWE-bench (Jimenez et al., 2023; Yang et al., 2024b) has motivated a range of agentic pipelines for software engineering tasks. Among them, SWE-agent (Yang et al., 2024a) and OpenHands (Wang et al., 2025a) are widely adopted frameworks that equip agents with tools for interacting with computing environments (terminal in Linux). Workflow-based methods such as Agentless (Xia et al., 2024) decomposes issue resolution into localization, repair, and validation subproblems. PatchPilot (Li et al., 2025) adds reproduction and refinement process in the workflow. CORTEXA (Sohrabizadeh et al., ) reflects the repository into an embedding space to enable efficient retrieval. Chen et al. (2025) builds the repository as a graph and applied graph-level searching tools for localization, and Wang et al. (2025a) furthermore integrated commit history as agent memory. RepoLens (Wang et al., 2025b) equip conceptual information of the repository to enable repo-level understanding. These pipelines are training-free, and yield competitive results.
To enable task-specific training, DeepSWE (Luo et al., 2025) and SWE-Swiss (He et al., 2025) employ reinforcement learning and achieve strong performance. SWE-Gym (Pan et al., 2024) builds a virtual environment to train an agent for this task. However, end-to-end training remains costly because patch evaluation requires executing Docker containers across numerous repositories, and the Docker container hinders guided search methods due to its non-serializability (Zainullina et al., 2025). Consequently, issue localization has emerged as a computationally efficient alternative, aiming to identify faulty components—at file or function level—rather than generating full patches.
@@ 3.3 Jump: Symbol Resolution
For a name symbol xx, candidate definitions are searched along a scope chain:
@@
For attribute symbols, it computes a (possibly union-valued) type tt for the receiver expression aa using type annotations, assignment flow analysis, function return types, and stub files (.pyi). Member resolution is then defined as:
@@
3.4 Reasoning: Action Loop
@@ 3.5 Reinforcement Learning
Table 1: Comparison of different agent pipelines on function-level and file-level metrics on SWE-bench_Verified. We use Qwen2.5-Instruct series as our base model. Bold numbers denote the best performance among same-size models; underline numbers denote the best training-free performance among same-size models; yellow background illustrates training-free RepoNavigator; blue background illustrates RepoNavigator trained with GRPO.
@@
closed-source Models
@@ 4.1 Experiment Setup
We extract valid samples from SWE-smith (Yang et al., 2025b) to form the training set. We apply Qwen2.5-7B-Instruct with RepoNavigator to sample each data for 16 times. A sample is abandoned if all 16 scores are zero. For validation, we test our method on SWE-bench_Verified (Jimenez et al., 2023), which is a human-verified subset of SWE-bench. We additionally test our method on a subset of SWE-bench_Pro (Yang et al., 2025b) (which is a new and more difficult benchmark) for generalization. For ground-truth locations, we directly use the locations in golden patches. All datasets are open-source and are built on real-world Github issues. We provide the strict definition of metrics in Appendix. B.
@@
Agentless Training Free 12.82 6.94 8.05 5.73 39.41 13.15 18.89 12.35
CoSIL Training Free 8.64 3.33 4.58 2.87 26.64 8.47 12.11 7.70
@@
Agentless Training Free 10.49 6.75 7.41 5.28 41.42 13.42 19.02 12.37
CoSIL Training Free 10.73 4.67 5.96 3.94 34.31 9.97 14.81 9.30
@@
Agentless Training Free 11.08 7.31 7.98 5.80 43.07 13.89 20.07 13.11
CoSIL Training Free 14.03 6.00 7.67 4.92 48.87 14.03 20.95 13.27
@@
RepoSearcher Training Free 2.52 2.46 2.31 1.75 9.00 2.52 3.81 2.41
@@
For the 7B model, we conduct GRPO with 8 NVIDIA Tesla-A100-80G GPUs. For the 14B and 32B model, we train it with 16 NVIDIA Tesla-A100-80G GPUs. We apply verl (Shen, 2024) as the training framework, and we apply vLLM (Kwon et al., 2023) as the inference engine. We train the model for 1 epoch, while the training batch size is fixed to 128 on 4k training samples filtered from SWE-smith, with maximum prompt length and max response length both set to 10240. Additionally, we rollout 8 times for each sample, and the temperature is set to 1.0 to encourage exploration. We use greedy decoding in the inference stage to ensure stable performance. More experimental details are provided in Appendix. B.
@@ 4.2 Effectiveness
We compare our method against LocAgent (Chen et al., 2025), CoSIL (Jiang et al., 2025), Agentless (Xia et al., 2024), OrcaLoca (Yu et al., 2025b), and RepoSearcher (Ma et al., 2025). LocAgent is distilled from Claude3.5-Sonnet, and RepoSearcher is distilled from Claude3.7-Sonnet. Detailed explanation of baseline methods are presented in Appendix. A.
@@ Results
As illustrated in Table. 1, on balanced metrics (Sample-F1 and IoU) for both function-level and file-level localization, our method surpasses all baseline methods with the same model size. Moreover, if we train RepoNavigator with GRPO, the 7B model surpasses 14B baselines, and our 14B model surpasses 32B baselines on Sample-F1 and IoU. This contributes to the validness of RepoNavigator furthermore. For 14B and 32B models, RepoNavigator achieves SOTA among all training-free methods. This implies that the tool we implement is effective and promising, and our single tool pipeline is better than previous multi-tools pipelines.
@@
To assess the generalizability of RepoNavigator, we present its performance on Python samples from the SWE-bench_Pro dataset (Yang et al., 2025b) in Table 2. The results on this dataset are consistent with those observed on SWE-bench_Verified. While we cannot fully exclude the potential influence of data leakage in SWE-bench_Verified, we can make a stronger claim regarding SWE-bench_Pro, as it was released after the publication of the Qwen2.5 series.
Figure 3: Training strategies comparison: comparison between RepoNavigator with training free, RFT, GRPO with pure outcome and hybrid reward on Qwen2.5-7B-Instruct.
@@ 4.3 Training Strategy Comparison
We investigate the capabilities of GRPO and our proposed hybrid reward function within the context of agentic training, using SFT as a primary benchmark. Experimental results presented in Fig. 3 indicate that our method significantly surpasses both SFT and GRPO with a vanilla reward function, validating the critical role of RL and hybrid reward mechanisms in enhancing agentic capabilities. While (Ma et al., 2025) emphasizes the necessity of an initial SFT phase, our ablation study (detailed in Appendix C) reveals the potential of training fully autonomous agents directly through RL. This observation challenges the requirement for a “cold start” and resonates with a growing body of literature advocating for RL-centric training paradigms (Liu et al., 2024a; Kirk et al., 2023; Jin et al., 2025b).